OpenStudy (anonymous):

integral t^2+1/t^3 dt where b=e^2 and a=e I tried using u-substitution but I couldn't figure it out , can someone help me start the problem ?

OpenStudy (luigi0210):

What did you use a u-sub?

OpenStudy (luigi0210):

*as a

OpenStudy (anonymous):

I tried with both, but I think u should be t^2+1

OpenStudy (luigi0210):

It should actually be u=t^3.

OpenStudy (anonymous):

so if I let u=t^3 then du/3t^2= dt ?

OpenStudy (anonymous):

oh heck no

OpenStudy (anonymous):

break it in to two pieces and you will be done quickly

OpenStudy (luigi0210):

Wait... hold on.

OpenStudy (luigi0210):

Right, sorry, not thinking straight right now.

OpenStudy (anonymous):

Okay so t^2/t^3 +1/t^3 and let u=t^3 ?

OpenStudy (luigi0210):

You could just integrate without a u-sub.

OpenStudy (anonymous):

really break it apart

OpenStudy (anonymous):

\[\frac{t^2+1}{t^3}=\frac{1}{t}+\frac{1}{t^3}\]

OpenStudy (anonymous):

Okay so when I broke it apart and used the du/3t^2 I get 1/3t^3 + 1/3t^2(t^3)

OpenStudy (anonymous):

stop with the u

OpenStudy (anonymous):

\(\frac{1}{t^3}=t^{-3}\)